OpenStudy (smurfy14):

Find the derivative. Please help!! f(x)=-3x[(-6)/(x+2)]

OpenStudy (callisto):

\[f(x) = -3x(\frac{-6}{x+2})\] Is this f(x) you've typed?

OpenStudy (smurfy14):

there should be a minus sign after -3x. sorry

OpenStudy (callisto):

\[f(x) = -3x - \frac{-6}{x+2}\] This?

OpenStudy (smurfy14):

ya

OpenStudy (callisto):

\[f(x) = -3x - \frac{-6}{x+2}\]\[f'(x) = \frac{d}{dx}(-3x) - \frac{d}{dx} (\frac{-6}{x+2})\]\[f'(x) = -3 - (-6)\frac{d}{dx} (\frac{1}{x+2})\]\[f'(x) = -3 +(6)\frac{d}{dx} (\frac{1}{x+2})\]\[f'(x) = -3 +(6)(-1) (\frac{1}{(x+2)^2})\]\[f'(x) = -3 -6 (\frac{1}{(x+2)^2})\]\[f'(x) = -3 - \frac{6}{(x+2)^2}\]

OpenStudy (smurfy14):

what does d/dx mean?

OpenStudy (callisto):

Differentiate (something) with respect to x ...

OpenStudy (smurfy14):

i dont know how you got rid of the x from -3x between lines 2 and 3

OpenStudy (callisto):

\[\frac{d}{dx} ax^n = anx^{n-1}\] Now, n =1, a=-3 \[\frac{d}{dx} -3x^1 = (-3)(1)x^{1-1} = -3\] Note that \[x^0 =1\]

OpenStudy (smurfy14):

oh ok i get it! one more questuon, why isnt the (x+2)^2 under the -3 as well?

OpenStudy (callisto):

They are two terms...

OpenStudy (smurfy14):

ok i got it thanks a bunch!!

OpenStudy (callisto):

Welcome!